Determine the volume when a 2.00-L sample of N2(g) is heated from 50.0 degrees Celsius to 300.0 degrees Celsius.
Answer #1

The relation we are wing Ti=50°C = (273+50) K = 323k V=2.00L T2 = 300° C = (273+300) k = 573k V ? Putting values = (2x 573) 3